Four years ago, ronnie was 6 times as old as his son. Six years ago, his age was 2 years more than eight times his son's age. how old are they now?

Answer by Aldorozos(172)

R = Ron's age today
S = Son's age today
Four year ago
R -4 = 6(S-4)
Six years ago
R-6 = 8(S-6) + 2
Now we have two equations and two unknowns
R -4 = 6(S-4)
R-6 = 8(S-6) + 2
One way to solve this is to subtract the second equation from the first equation. This eliminates R and we end up with one equation in terms of S
R-6 -(R-4) = 8(S-6) + 2 - 6(S-4)
-2 = 2S -22
20 = 2S
and therefore S = 20/2 = 10
Now that we have the age of the son which is 10 years old now, we can put this number in the first equation and calculate the age of the father which is R
R-4 = 6(10-4)
R-4 = 36
R= 40 Therefore the father is 40 years old.